# Implicit geometric regularization in flow matching via density weighted Stein operators

### v1

@@ 4.2 Latent-flow modelling of CIFAR-10
Thus, minimizing $\mathrm{GSC}$ creates a “geometric anchor” that prevents the model from overfitting to the sparse ambient space. Future work may explore scalable approximations of the rigorous penalty, such as diagonal approximations or last-layer Laplace approximations, to bridge the gap between the rigorous theory and deep learning practice.
@@
Figure 3: Analysis of geometric regularization and hyperparameter selection. (a) Micro-level analysis: The distribution of the Jacobian norm $\|\nabla_{x}v_{\theta}\|_{F}$ (roughness) for the baseline ($\gamma=0$) and our method ($\gamma=1.0$). Our method suppresses high-frequency oscillations, resulting in a smoother vector field (concentrated at lower values). (b) Macro-level selection: The GSC score across different $\gamma$ values. The GSC, computed as a sum of normalized bias (Inlier MMD) and variance penalty (Smoothness), identifies $\gamma=1.0$ as the optimal trade-off.
@@ 4.3 Stress test under contaminated latents
Although our main experiments use clean data and focus on geometric regularization on the data manifold, it is natural to ask whether the same density-weighted geometry also confers classical robustness in the presence of explicit contamination. To probe this aspect, we construct a synthetic contaminated latent dataset by injecting a small fraction of adversarial latents into the training set. These latents are sampled from a broader Gaussian distribution that lies far from the main data manifold, but spatially proximate enough to distract the learning process. Figure 4 visualizes the generated samples projected onto the first two principal components. Standard Flow Matching ($\gamma=0$, Left) fails to distinguish between inliers and outliers: the learned flow attempts to fit all target signals, treating noise as valid data. This results in distorted manifolds and poor generalization. In contrast, $\gamma$-Flow Matching ($\gamma=0.5$, Right) downweights the adversarial latents via $\hat{p}_{t}^{\gamma}$, effectively ignoring their contribution. The learned flow remains aligned with the main data manifold, yielding robust samples that are visually indistinguishable from the clean-data case.
Figure 4: Stress test under contaminated latents. Left: Standard FM ($\gamma=0$) is misled by adversarial latents and learns a distorted manifold. Right: $\gamma$-FM ($\gamma=0.5$) downweights the contaminants via $\hat{p}_{t}^{\gamma}$ and preserves the structure of the inlier manifold. This experiment illustrates that the same density-weighted geometry used to organize learning on the data manifold also yields classical robustness to explicit contamination.
@@
Finally, we examine the relationship between vector field smoothness and the number of function evaluations (NFE) required by an ODE solver. We consider a range of NFE values and measure the Fréchet distance between generated images and real images in the latent space. With a smooth vector field (achieved by $\gamma$-FM), a coarse solver with low NFE already yields high-quality samples, since the local truncation errors do not accumulate catastrophically. In contrast, with a rough vector field (standard FM), reducing the step size (increasing NFE) does not necessarily improve sample quality: the solver is approximating a poor flow that overfits the voids. High NFE efficiency is often associated with flow rectification methods that enforce straight trajectories (liu2023flow). Our results suggest that $\gamma$-weighting achieves a similar efficiency gain by smoothing the vector field in void regions, effectively removing the stiff components of the dynamics.
Figure 5: Efficiency Analysis (NFE vs. quality). Comparison of generation quality across different ODE solver steps. Red ($\gamma$-FM): Achieves optimal quality at relatively low NFE, reflecting a straight and well-conditioned flow. Blue (Standard FM): Suffers from an Inverse Precision Paradox, where increasing solver precision does not correct for a vector field that has learned the wrong geometry.
@@ 5 Conclusion
We introduced $\gamma$-Flow Matching as a robust framework for continuous normalizing flows. Our analysis establishes that the $\gamma$-weighted objective serves as an implicit geometric regularizer, yielding smoother vector fields by filtering out chaotic signals in low-density regions. Theoretical connections to the porous medium equation provide a physical grounding for this behavior, interpreting void rejection as a finite-propagation phenomenon. Experiments confirm that $\gamma$-FM significantly reduces vector field roughness and improves sampling efficiency without compromising generation quality. This work suggests that density-weighted regression should be considered a standard tool for high-dimensional generative modeling, where classical robustness emerges naturally from the underlying geometry.
@@ Appendix A $\gamma$-Stein connections and geodesic viewpoint
The $\gamma$-divergence $D_{\gamma}(p,q)$ generates a Riemannian metric $g^{(\gamma)}$ and a pair of dual affine connections $(\nabla^{(\gamma)},\nabla^{(\gamma)*})$ on the statistical manifold $\mathcal{M}=\{q_{\theta}\}$ according to the theory of minimum contrast geometry (Eguchi1992; eguchi2009information). Ignoring an irrelevant constant factor, the induced metric takes the form
@@
Following the general construction of Eguchi1992, the affine connection $\nabla^{(\gamma)}$ can be realized as the $L^{2}$-projection of the $\theta$-derivative of score fields back onto the score span. More precisely, for each coordinate direction $\partial_{k}$ we consider $\partial_{k}s_{i}=\partial_{\theta^{k}}\partial_{\theta^{i}}\log q_{\theta}$ as an element of $\mathcal{H}_{\theta,\gamma}$ and define the projection
@@
that is, $u_{t}$ is covariantly constant under the $\gamma$-Stein connection along the curve. This provides a geometric counterpart to the sample-space flow governed by the nonlinear Fokker–Planck equation, and suggests a dual picture in which $\gamma$-flow matching approximately follows geodesics on the statistical manifold endowed with the $\gamma$-Fisher metric and the $\gamma$-Stein connection. A more detailed study of this escort geometry is left for future work; see, e.g., Ohara2010; Matsuzoe2017 for related developments on escort distributions.
Appendix B Hyperparameter Sensitivity
A potential concern with density-weighted regression is the computational overhead introduced by the $k$-Nearest Neighbors ($k$-NN) estimation within each training batch. To address this, we conducted an ablation study measuring the wall-clock time per iteration and training stability across varying numbers of neighbors $k\in\{5,10,20,50,100\}$.
Experimental Setup. We used a batch size of $B=512$ and measured the forward-backward pass time on a single NVIDIA T4 GPU. The results are averaged over 1000 iterations.
Results. As shown in Table 2, the computational cost is effectively invariant to $k$. The average time per iteration remains approximately 2.0–2.5 ms regardless of $k$. This is because the computational complexity is dominated by the pairwise distance calculation ($O(B^{2})$), which is highly parallelized on GPUs, while the subsequent top-$k$ selection adds negligible overhead for typical batch sizes. Furthermore, the loss values indicate that the training stability is robust to the choice of $k$. Thus, $\gamma$-FM adds minimal computational burden compared to standard Flow Matching.
Table 2: Ablation study on neighbor size $k$. Time per iteration is nearly constant.
@@ Appendix B Hyperparameter Sensitivity
Appendix C Latent-flow modelling of CIFAR-10
This appendix provides implementation details and additional visual results for the latent-flow CIFAR-10 experiments in Section 4.2. To validate the quality of the latent representation, we visually inspect the reconstruction capabilities of the pre-trained autoencoder. Figure 6(a) displays the reconstruction results using an Exponential Moving Average (EMA) of the weights. The images are arranged in an interleaved manner, where each original input image is immediately followed by its corresponding reconstruction. The results indicate that the autoencoder preserves sufficient structural and semantic details to serve as a basis for the generative model.
Subsequently, we trained the flow matching model within this fixed latent space. Figure LABEL:fig:gamma_samples shows the generated samples obtained from the model with $\gamma=0.0$. The decoding process utilized a retrieval-skip strategy with parameters $k=2$ and $\tau=0.5$. These samples demonstrate the model’s ability to synthesize coherent images from the learned latent distribution.
(a) AE reconstructions (CIFAR-10).
(b) $\gamma=0.0$ (retrieval-skip: $k=2$, $\tau=0.5$).
(c) $\gamma=0.5$ (retrieval-skip: $k=2$, $\tau=0.5$).
(d) $\gamma=1.0$ (retrieval-skip: $k=2$, $\tau=0.5$).
Figure 6: Autoencoder reconstruction quality and random samples from latent flow matching under different $\gamma$ values (CIFAR-10).

### v2

@@ 4.3 Evaluation on latent CIFAR-10
Thus, the GSC may provide a useful geometric diagnostic for increasingly conservative low-density regularization, even though in the present experiments it is not used as a stand-alone selection rule. Future work may explore scalable approximations of the rigorous penalty, such as diagonal approximations or last-layer Laplace approximations, to bridge the gap between the rigorous theory and deep learning practice.
@@
Figure 4: Auxiliary diagnostics for smoothness and $\gamma$ selection. (a) Micro-level analysis: Illustrative roughness histogram for the baseline ($\gamma=0$) and our method ($\gamma=1.0$). The weighted objective is associated with a milder roughness profile and fewer extreme oscillatory regions, although the two distributions still overlap substantially. (b) Macro-level selection: The GSC score across different $\gamma$ values. In the shared-time protocol, this curve is interpreted jointly with validation FID and decoded-image diagnostics rather than as a stand-alone optimum certificate.
Practical stability of the shared-time proxy.
A natural concern is whether the shared-time $k$-NN weighting proxy becomes unreliable when the minibatch is small. To address this point directly, we report an additional minibatch-size ablation in Appendix D.1, where the latent CIFAR-10 experiment is repeated under the same frozen autoencoder, decoding pipeline, optimizer, and training budget while varying only the minibatch size. The main finding is that, in the tested latent-CIFAR regime, the proxy retains a stable nontrivial dynamic range across the tested batch sizes rather than collapsing at smaller minibatches. At the same time, both FM and $\gamma$-FM improve as the batch size increases, indicating that the degradation at small $B$ is not specific to the density-weighted objective. In particular, Appendix D.1 shows that the final ESS/B/B stays near $0.52$ across the tested batch sizes, indicating that the proxy retains a stable nontrivial dynamic range rather than degenerating into nearly uniform or highly singular weights. We therefore interpret the shared-time surrogate as a practically stable within-batch ordering device in the regime studied here, rather than as a general guarantee for high-dimensional minibatch density estimation. Moreover, the same ablation shows that at low to moderate batch sizes, $\gamma$-FM tends to produce smoother learned vector fields than FM while maintaining broadly comparable recall and coverage.
@@ 4.4 Stress test under contaminated latents
While the density-weighted objective is well defined under both clean and contaminated targets, the clean-target rerun suggests that its empirical benefit is most visible when the regression geometry is distorted by out-of-support or contaminated latents. We therefore retain a more adversarial latent-contamination stress test to examine that regime directly.
These latents are sampled from a broader Gaussian distribution that lies away from the main data manifold. Figure 5 visualizes the generated samples projected onto the first two principal components. Standard Flow Matching ($\gamma=0$, Left) is more easily distracted by these contaminants: the learned flow attempts to fit both inlier and adversarial target signals, resulting in a distorted latent structure. In contrast, $\gamma$-Flow Matching ($\gamma=0.5$, Right) downweights the contaminants via $\hat{p}_{t}^{\gamma}$, thereby preserving closer alignment with the dominant inlier manifold. This stress test therefore illustrates the regime in which the proposed weighting is most helpful: not necessarily already-clean targets, but target geometries that are distorted by contamination or poor support.
Figure 5: Stress test under a more adversarial latent-contamination pattern. Left: Standard FM ($\gamma=0$) is more easily misled by adversarial latents and learns a distorted manifold. Right: $\gamma$-FM ($\gamma=0.5$) downweights the contaminants via $\hat{p}_{t}^{\gamma}$ and better preserves the dominant inlier structure. This figure is intended to highlight the regime in which density weighting is most useful. In a separate clean-target rerun, where the target geometry was not distorted by contamination, the same weighting mechanism remained active but did not yield a comparably clear gain in decoded-image quality.
@@
Finally, we examine the relationship between vector field smoothness and the number of function evaluations (NFE) required by an ODE solver. We consider a range of NFE values and measure the Fréchet distance between generated images and real images in the latent space. In our experiments, smoother learned vector fields were often associated with better quality at lower NFE. This should be interpreted as an empirical tendency in the tested setting rather than as a universal consequence of the method. In contrast, with a rough vector field (standard FM), reducing the step size (increasing NFE) does not necessarily improve sample quality: the solver is approximating a poor flow that overfits the voids. High NFE efficiency is often associated with flow rectification methods that enforce straight trajectories (Liu et al., 2023). Our results suggest that $\gamma$-weighting achieves a similar efficiency gain by smoothing the vector field in void regions, effectively removing the stiff components of the dynamics.
Figure 6: Efficiency Analysis (NFE vs. quality). Comparison of generation quality across different ODE solver steps in the tested latent-CIFAR setting. Red ($\gamma$-FM): Reaches near-optimal quality at relatively low NFE, consistent with a smoother and better-conditioned flow in this experiment. Blue (Standard FM): Shows diminishing returns as NFE increases, suggesting that finer ODE integration does not compensate for a less well-conditioned vector field–a tendency we refer to informally as an Inverse Precision Paradox. This behavior is an empirical observation in the present setting and should not be interpreted as a universal consequence of standard flow matching.
@@ 5 Conclusion
We introduced $\gamma$-Flow Matching as a density-weighted alternative to standard flow matching for high-dimensional generative flows. The presentation separates what is formally established from what is offered as geometric or spectral intuition: the weighted objective suppresses low-density variance contributions, the porous-medium analogy clarifies a “void rejection” mechanism, and the weighted Dirichlet viewpoint suggests a route toward smoother learned vector fields. Empirically, the evaluation protocol is designed to test these claims under a shared-time density-estimation scheme, with explicit comparison to FM++Jac and with decoded-image metrics that quantify both fidelity and diversity.
@@
As with many efficiency-improving advances in generative modeling, the proposed method may support beneficial uses such as simulation and data augmentation, but it may also lower the cost of producing synthetic content at scale. The present paper is intended as a methodological contribution, and we do not claim application-specific safeguards beyond standard responsible use and evaluation in controlled research settings.
Code availability
The Python notebooks used to reproduce the simulation studies are available at https://github.com/shinto-eguchi/gamma-flow-matching.
Acknowledgements
The author is grateful to the Action Editor and the anonymous reviewers for their careful reading, constructive comments, and encouraging assessment of the work. Their feedback helped improve the clarity, positioning, and presentation of the paper. The author also thanks the TMLR editorial team for coordinating the review process.
@@ Appendix A $\gamma$-Stein connections and geodesic viewpoint
The $\gamma$-divergence $D_{\gamma}(p,q)$ generates a Riemannian metric $g^{(\gamma)}$ and a pair of dual affine connections $(\nabla^{(\gamma)},\nabla^{(\gamma)*})$ on the statistical manifold $\mathcal{M}=\{q_{\theta}\}$ according to the theory of minimum contrast geometry (Eguchi, 1992, 2009). Ignoring an irrelevant constant factor, the induced metric takes the form
@@
Following the general construction of Eguchi (1992), the affine connection $\nabla^{(\gamma)}$ can be realized as the $L^{2}$-projection of the $\theta$-derivative of score fields back onto the score span. More precisely, for each coordinate direction $\partial_{k}$ we consider $\partial_{k}s_{i}=\partial_{\theta^{k}}\partial_{\theta^{i}}\log q_{\theta}$ as an element of $\mathcal{H}_{\theta,\gamma}$ and define the projection
@@
that is, $u_{t}$ is covariantly constant under the $\gamma$-Stein connection along the curve. This provides a geometric counterpart to the sample-space flow governed by the nonlinear Fokker–Planck equation, and suggests a dual picture in which $\gamma$-flow matching approximately follows geodesics on the statistical manifold endowed with the $\gamma$-Fisher metric and the $\gamma$-Stein connection. A more detailed study of this escort geometry is left for future work; see, e.g., Eguchi and Kato (2010); Matsuzoe (2017) for related developments on escort distributions.
Appendix B Hyperparameter and proxy sensitivity
A potential concern with density-weighted regression is the computational overhead introduced by the $k$-Nearest Neighbors ($k$-NN) estimation within each training batch. To address this, we first evaluate the computational cost and stability of the practical $k$-NN surrogate used in our implementation. We then examine how the surrogate behaves as the minibatch size $B$ and the ambient dimension $D$ vary under the shared-time protocol.
The purpose of this appendix is operational rather than asymptotic: we do not claim a universal high-dimensional guarantee for $k$-NN density estimation. Instead, we document how the within-batch geometry and the induced practical weights behave in the regime relevant to our experiments.
Appendix C Computational cost of the $k$-NN surrogate
We first measure the wall-clock time per training iteration while varying the number of neighbors $k\in\{5,10,20,50,100\}$. All experiments use batch size $B=512$ and are executed on a single NVIDIA T4 GPU. The reported values are averaged over 1000 iterations.
Results. Table 2 shows that the computational cost is effectively invariant to the choice of $k$. The average time per iteration remains approximately 2.02.02.52.5 ms across all tested values. This occurs because the dominant cost arises from the pairwise distance computation ($O(B^{2})$), which is highly parallelized on modern GPUs. The additional top-$k$ selection step introduces negligible overhead for typical batch sizes. The loss values also show that training stability is insensitive to the precise choice of $k$. Therefore, the practical surrogate adds minimal computational burden compared with standard Flow Matching.
Table 2: Ablation study on neighbor size $k$. Time per iteration remains nearly constant.
@@ Appendix C Computational cost of the $k$-NN surrogate
C.1 Sensitivity to batch size and ambient dimension
To better understand the geometric behavior of the practical surrogate, we conducted a diagnostic experiment varying both the minibatch size $B$ and the effective ambient dimension $D$ under the shared-time protocol.
The analysis focuses on three quantities:
  • the $k$-NN distance statistic $d_{k}(x)$,
  • the inverse-distance proxy $\rho_{k}(x)=1/(d_{k}(x)+\varepsilon)$,
  • the induced practical weights $w(x)$ used in the $\gamma$-weighted regression objective.
Dimension sensitivity. At fixed batch size $B=512$, the raw $k$-NN geometry becomes increasingly concentrated as the ambient dimension increases. The coefficient of variation of the mean-$k$ distance decreases from $\mathrm{CV}(d_{k})=0.087$ at $D=32$ to $0.044$ at $D=256$. Similarly, the contrast ratio $q_{0.8}(d_{k})/q_{0.2}(d_{k})$ decreases from $1.158$ to $1.078$. The same trend appears for the proxy $\rho_{k}$, where $\mathrm{CV}(\rho_{k})$ decreases from $0.086$ to $0.045$. These results are consistent with the well-known concentration of pairwise distances in higher dimensions.
Batch-size sensitivity. At fixed dimension $D=256$, increasing the minibatch size from $B=128$ to $B=1024$ slightly strengthens the within-batch ordering signal. The coefficient of variation $\mathrm{CV}(d_{k})$ increases from $0.041$ to $0.048$, and the contrast ratio $q_{0.8}(d_{k})/q_{0.2}(d_{k})$ increases from $1.071$ to $1.085$. Thus, larger batches provide a modestly sharper local ordering signal.
Behavior of the practical weights. Importantly, the final weights are obtained through a rank-based monotone transform of the proxy rather than from its absolute scale. Consequently, summary statistics of the induced weights, such as $\mathrm{ESS}(w)/B$, remain nearly invariant to the ambient dimension. This behavior is by design: the practical surrogate is intended to provide a stable within-batch ordering signal rather than to estimate densities on an absolute scale. In practice, the surrogate preserves sufficient ordering information even though the raw $k$-NN geometry becomes less contrasted in higher dimensions.
Table 3: Sensitivity of the $k$-NN geometry and the induced practical weights to batch size $B$ and ambient dimension $D$.
Setting $B$ $D$ $\mathrm{CV}(d_{k})$ $q_{0.8}(d_{k})/q_{0.2}(d_{k})$ $\mathrm{CV}(\rho_{k})$ $\mathrm{ESS}(w)/B$
Fixed $D=256$ (varying $B$)
128 256 0.0407 1.071 0.0412 0.603
256 256 0.0414 1.073 0.0421 0.603
512 256 0.0443 1.078 0.0454 0.603
1024 256 0.0481 1.085 0.0499 0.603
Fixed $B=512$ (varying $D$)
512 32 0.0873 1.158 0.0864 0.603
512 64 0.0668 1.119 0.0671 0.603
512 128 0.0515 1.091 0.0522 0.603
512 256 0.0445 1.078 0.0455 0.603
Figure 7 visualizes these trends. The results show that although the raw geometric contrast weakens as dimension increases, the rank-based practical surrogate remains operationally stable for the range of $B$ and $D$ used in our experiments.
Figure 7: Operational sensitivity of the practical $k$-NN surrogate. Top: behavior of the raw $k$-NN distance statistic. Bottom: behavior of the induced practical weights and proxy.
Appendix D Latent-flow modelling of CIFAR-10
Implementation details.
The latent CIFAR-10 experiments use a robust-target setup with $n_{\mathrm{train}}=10000$ images and contamination fraction $0.40$, consisting of $6000$ inlier-class images and $4000$ outlier-class images. All latent-space experiments use the same frozen autoencoder and the same decoding pipeline across FM, shared-time $\gamma$-FM, and FM++Jac. The autoencoder is a pre-trained EMA checkpoint of a CleanAE architecture with base width $64$ and latent dimension $d_{z}=256$.
The latent flow backbone is a VelocityMLP with hidden width $512$ and depth $3$, trained for $60$ epochs with batch size $512$ using Adam with learning rate $10^{-3}$ and EMA decay $0.999$. In the shared-time protocol, each minibatch draws a single $t\sim\mathcal{U}[0,1]$ and constructs all particles at that common time. The practical weighting uses a rank-based monotone $k$-NN surrogate with $k=10$, lower and upper rank quantiles $0.20$ and $0.80$, rank scale $2.0$, and weight clipping to $10^{-3},5\times 10^{1}]$.
For FM++Jac, the penalty coefficient is $\lambda_{J}=10^{-3}$, and the Jacobian norm is estimated by a single Hutchinson probe with Rademacher noise during both training and evaluation. Sampling uses a fixed-step midpoint integrator with $100$ steps. Decoding uses retrieval-skip with a skip bank of $600$ encoded training images, $k=2$, and $\tau=0.5$. Latent MMD is computed with subsample size $2000$, and smoothness is evaluated with $2000$ random latent samples and one Rademacher probe. FID, Recall, and Coverage are computed from $5000$ generated images, with metric batch size $128$ and coverage parameter $k=5$; the reference set consists of the $1000$ test images of the inlier class. Multi-seed runs use seeds $42$, $1042$, and $2042$ on a single NVIDIA Tesla T4 GPU.
This appendix provides additional visual results for the latent-flow CIFAR-10 experiments in Section 4.2. We make the experimental protocol explicit by reporting the autoencoder architecture, latent dimensionality, flow backbone, shared-time minibatch construction, practical $k$-NN surrogate parameters, Jacobian-penalty implementation, fixed-step sampling protocol, decoding protocol, the number of generated samples used for FID/Recall/Coverage, random seeds, and hardware.
Subsequently, we trained the flow matching models within this fixed latent space. Figure 8 shows representative decoded samples obtained from FM, $\gamma$-FM (shared-$t$), and FM++Jac. The decoding process uses retrieval-skip with parameters $k=2$ and $\tau=0.5$. In the comparison, the same frozen decoder and the same retrieval-skip settings are used for FM, $\gamma$-FM (shared-$t$), and FM++Jac so that image-space differences reflect changes in the latent-flow objective rather than changes in the representation or decoding pipeline. These samples are intended as a qualitative complement to the quantitative metrics, showing that the overall decoded sample quality remains reasonable across the compared methods under a fixed latent representation.
D.1 Minibatch-size sensitivity of the shared-time weighting proxy
A natural concern is whether the shared-time $k$-NN proxy used in $\gamma$-FM becomes unreliable when the minibatch is small. To examine this point, we repeated the latent CIFAR-10 experiment under the same latent representation, frozen autoencoder, decoding pipeline, optimizer, and training budget as in the main experiment, varying only the minibatch size
$B\in\{128,256,512,1024\}.$
We compared shared-time FM and shared-time $\gamma$-FM with $\gamma=1$, and report means and standard deviations over three random seeds in Table 4.
The main observation is that the weighting proxy did not collapse at smaller batch sizes in this regime. For $\gamma$-FM, the final effective sample size ratio $\mathrm{ESS}/B$ remained between $0.512$ and $0.526$ across all tested values of $B$, while the final weight quantile ratio $q_{0.95}/q_{0.05}$ stayed near $20.09$ throughout. Thus, the proxy retained a stable, nontrivial dynamic range rather than degenerating as $B$ decreased.
As expected, image-space quality improved with minibatch size for both FM and $\gamma$-FM. Hence, the degradation at small $B$ should not be interpreted as a pathology specific to density weighting; rather, it reflects the general difficulty of minibatch training in this contaminated latent-generation setting. Within this shared trend, the most consistent advantage of $\gamma$-FM appears in the geometry of the learned field: at low to moderate batch sizes ($B=128,256,512$), the smoothness proxy is lower for $\gamma$-FM than for FM, while recall and coverage remain broadly comparable. This is consistent with the intended role of the weighting scheme as an implicit geometric regularizer.
As a complementary sanity check, we also repeated the same minibatch-size experiment under a clean-target latent protocol, that is, without outlier contamination in the training targets. In this rerun, the weighting mechanism remained clearly active, with final $\mathrm{ESS}/B$ near $0.52$ and the final weight quantile ratio near $20$ across batch sizes. However, the decoded-image advantage largely disappeared. At $B=1024$, for example, shared-time FM achieved slightly better FID, Recall, and Coverage than shared-time $\gamma$-FM, while the smoothness values became nearly identical. This clean-target rerun is therefore included mainly as a sanity check on the mechanism rather than as a regime in which $\gamma$-FM is expected to dominate FM.
(a) FM (shared-$t$) samples (retrieval-skip: $k=2$, $\tau=0.5$).
(b) $\gamma$-FM with $\gamma=0.5$ (retrieval-skip: $k=2$, $\tau=0.5$).
(c) $\gamma$-FM with $\gamma=1.0$ (retrieval-skip: $k=2$, $\tau=0.5$).
(d) FM++Jac with $\lambda_{J}=10^{-3}$ (retrieval-skip: $k=2$, $\tau=0.5$).
Figure 8: Representative decoded samples for FM (shared-$t$), $\gamma$-FM ($\gamma=0.5,1.0$), and FM++Jac under the latent CIFAR-10 protocol. All methods use the same frozen decoder and the same retrieval-skip settings ($k=2$, $\tau=0.5$). This figure is intended as a qualitative complement to the quantitative comparisons in Table 1.
Table 4: Minibatch-size ablation for shared-time FM and shared-time $\gamma$-FM ($\gamma=1$) on latent CIFAR-10 under the robust-target protocol. Each entry is the mean $\pm$ standard deviation over three random seeds. Lower FID and smoothness are better, while higher Recall and Coverage are better. For FM, $\mathrm{ESS}/B=1$ by construction because the weights are uniform.
[…]
FM (shared-$t$) 128 $64.321\pm 0.757$ $0.370\pm 0.012$ $0.893\pm 0.012$ $64.813\pm 3.335$ $1.000$ $1.00$
FM (shared-$t$) 256 $62.407\pm 2.303$ $0.402\pm 0.012$ $0.907\pm 0.005$ $33.768\pm 4.525$ $1.000$ $1.00$
FM (shared-$t$) 512 $59.063\pm 0.759$ $0.388\pm 0.008$ $0.915\pm 0.006$ $9.136\pm 0.110$ $1.000$ $1.00$
FM (shared-$t$) 1024 $53.054\pm 0.965$ $0.406\pm 0.020$ $0.938\pm 0.004$ $1.209\pm 0.175$ $1.000$ $1.00$
$\gamma$-FM ($\gamma=1$) 128 $66.076\pm 0.657$ $0.372\pm 0.010$ $0.889\pm 0.011$ $50.795\pm 2.236$ $0.512$ $20.09$
$\gamma$-FM ($\gamma=1$) 256 $64.339\pm 1.434$ $0.384\pm 0.017$ $0.877\pm 0.015$ $27.697\pm 2.384$ $0.518$ $20.09$
$\gamma$-FM ($\gamma=1$) 512 $61.198\pm 0.735$ $0.395\pm 0.009$ $0.915\pm 0.011$ $8.788\pm 0.194$ $0.524$ $20.09$
$\gamma$-FM ($\gamma=1$) 1024 $54.223\pm 0.486$ $0.412\pm 0.019$ $0.933\pm 0.005$ $1.253\pm 0.155$ $0.526$ $20.09$